Highest Modulus of the ring counter made up of 5 flip-flops is 25.State true or false, also give a reason for your answer.
1
Expert's answer
2018-03-24T07:55:07-0400
We know, N = 2^n
Where, N = Highest modulus of the ring counter
n = number of flip- flops
Now, here n = 5 , then N = 2^5 = 32

Answer: So, the statement is false.